OpenStudy (anonymous):

Jose has $2.50 in nickels and dimes in his jar. If he has 11 more nickels than dimes, how many dimes does he have?

OpenStudy (apoorvk):

no. of nickels=x no. of dimes=y so, x=y+11 and total value is $2.50, that is 250 cents so, 5x+10y=250 simplifies to x+2y=50 solve the two linear equations to get x and y, and enjoy!!

OpenStudy (anonymous):

so how many dimes does jose have>?

OpenStudy (apoorvk):

what do you think?? post and verify!

OpenStudy (anonymous):

13??

OpenStudy (apoorvk):

listen i am just confirming 'cause i don't live in the states.. nickel is 5cents and dime is 10.. right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it would be 15 correct?

OpenStudy (apoorvk):

nah.. 13 was right! (for dimes)

OpenStudy (anonymous):

oh Ok thanxx

OpenStudy (anonymous):

he sum of two numbers is 60. Their difference is 12. What is the smaller number?

OpenStudy (anonymous):

the*